OpenStudy (emmynimmy):

The equation y = 35.25x + 40 gives the labor cost y of repairing a car if it takes x hours to repair the car. Which statement is true? For every hour of labor, the cost decreases by $40. For every hour of labor, the cost decreases by $35.25. For every hour of labor, the cost increases by $40. For every hour of labor, the cost increases by $35.25.

OpenStudy (teddyiswatshecallsme):

I believe its D.

OpenStudy (emmynimmy):

Are you sure?

OpenStudy (emmynimmy):

The equation y = 4.75x + 20 gives the cost y of towing a car if the car is towed x miles. According to the model, what is the best estimate of the cost for towing a car 22 miles? $24.75 $46.25 $104.50 $124.50

OpenStudy (teddyiswatshecallsme):

Well its not negative so it wouldn't decline, and 40 doesn't represent the hourly cost so yeah. lol cx

OpenStudy (emmynimmy):

thanks! can you help with the other question?

OpenStudy (johnweldon1993):

\[\large y = 4.75x + 20\] where 'x' is the number of miles...so since we want to know the cost to go 22 miles we have \[\large y = 4.75(22) + 20\] \[\large y = 104.5 + 20\] \[\large y = 124.5\]

OpenStudy (teddyiswatshecallsme):

Yeah. Briefly. The answer for the second is 104.50

OpenStudy (teddyiswatshecallsme):

4.75(22) = 104.50

OpenStudy (johnweldon1993):

...close...but there is a constant added after that

OpenStudy (emmynimmy):

so that is not the answer?

OpenStudy (johnweldon1993):

*refers to my post above :)

OpenStudy (emmynimmy):

would it be 124?

OpenStudy (emmynimmy):

Awesome thanks!

OpenStudy (emmynimmy):

The correlation coefficient (r) between the number of miles driven x and the number of gallons of gasoline used y is 0.925. What percent of the variation in the number of gallons of gasoline can be explained by differences in the number of miles driven? 7.5% 14.4% 85.6% 92.5%

OpenStudy (teddyiswatshecallsme):

Oh i forgot about the 20.. Yeah, John's right. lolol. Most of these questions are easy you just plug in numbers for x or y and then just add or subtract or whatever.

OpenStudy (emmynimmy):

okay what about that ^^^

OpenStudy (johnweldon1993):

Oh didn't see a reply So the correlation coefficient r = 0.925 What we need for variability is R^2 ...so .925^2 = 0.8556 And multiply that by 100 to get the percent which is 85.6%

OpenStudy (emmynimmy):

Thank you so much!

OpenStudy (emmynimmy):

@johnweldon1993

OpenStudy (johnweldon1993):

Oh crap, I hate this one haha I remember the long tedious steps to do this So I'm gonna cheat...1 sec :P

OpenStudy (emmynimmy):

lol okay thanks!

OpenStudy (johnweldon1993):

First choice sorry it took so long excel didnt wanna open for me lol :P

OpenStudy (emmynimmy):

no worries! thankyou!!

OpenStudy (johnweldon1993):

np :)

OpenStudy (emmynimmy):

Sorry, I'm literally trying so hard to turn all of my work in on time. Math is my worst subject.

OpenStudy (emmynimmy):

@johnweldon1993